EXPERIMENTAL CHARACTERIZATION OF THE He + I^{35}Cl(E,v^{†}=11,12) AND He + I^{35}Cl(β,v^{†}=0-2) IN-TERMOLECULAR POTENTIAL ENERGY SURFACES

\underline{JOSHUA P. DARR} and RICHARD A. LOOMIS, Department of Chemistry, Washington University, One Brookings Drive, CB 1134, Saint Louis, MO 63130.

Two-photon excitation of the T-shaped and linear He⋯I^{35}Cl(X ^{1}Σ^{+},v′′=0) complexes are used to access multiple intermolecular vibrational levels within the He + I^{35}Cl(E0^{+} ^{3}P_{2},v^{†}=11,12) and He + I^{35}Cl(β1 ^{3}P_{2},v^{†}=0-2) ion-pair state intermolecular potentials. The excitation utilizes different metastable intermolecular levels within the He + I^{35}Cl(B 3Π_{0+},v′=2,3) and He + I^{35}Cl(A ^{3}Π_{1},v′=15) potentials, and thus varying Franck-Condon windows, to access levels with varying amounts of vibrational excitation within the ion-pair state intermolecular potential. The He + ICl(E,v^{†}=11,12) and He + ICl(β,v^{†}=0-2) intermolecular potentials are found to be nearly identical with an overall minimum in the near T-shaped orientation and binding energies D _{0}^{†} = 40 cm^{−1}. Intermolecular stretching and bending frequencies are measured to be 25 and 13 cm^{−1}, respectively. Since common intermolecular levels are accessed by transitions from both the T-shaped and linear He⋯ICl(X,v′′=0) ground state conformers, the relative binding energies of the conformers can be directly measured; the linear conformer is found to be 5.4(2) cm^{−1} more strongly bound than the T-shaped conformer.